OpenStudy (kendricklamar2014):

Simplify the following expression. Show each step. 33xy/3x <---- fraction

OpenStudy (kendricklamar2014):

I think its 11y

OpenStudy (gorv):

\[\frac{ 33*x*y }{ 3*x }=11*y\]

OpenStudy (gorv):

3 will cut 33 11 time and x get cancelled

OpenStudy (kendricklamar2014):

so the answer is 11y right

OpenStudy (kendricklamar2014):

@gorv

OpenStudy (gorv):

yeah

OpenStudy (kendricklamar2014):

thanks